the one by on step proof for a-(a-b)=2a-b

Answer by swincher4391(1107)

I think you may have a typo. I believe it's a-(-a+b) = 2a-b or maybe a-(b-a) = 2a-b.
Anyway,
a-(b-a) = 2a-b
Distribute
a - b + a = 2a -b
Combine like terms
2a - b = 2a -b

Or if its
a - (-a+b) = 2a -b
Distribute
a +a -b = 2a-b
Combine like terms
2a - b = 2a - b